What is the formula for the equation which has the the points (1,8) and (-1,12)?
Answer by Slayer(3)

(1,8) and (-1,12)
REMEMBER: The equation should look like " y=mx+b "
m = the slope
b = y intercept
First find the slope FORMULA: y2 - y1 / x2 - x1.
So; 12-8 / -1-1
= 4/-2 = -2
Fill in the slope; y=-2x+b; now all we need in the "b"
Take one of those points for instance (1,8) Fill them into the equation.
8=-2(1)+b
8=-2+b
+2 +2
10 = b
Now the eqiation is y = -2x + 10